HUBBLE SPACE TELESCOPE DAILY REPORT #1570 PERIOD COVERED: 0000Z (UTC) 02/13/96 - 0000Z (UTC) 02/14/96 Daily Status Report as of 045/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5971 (HST Medium Deep Survey) The WF/PC-2 was used to make a long exposure medium deep sky survey = observations at latitudes larger than 10 degrees taken in parallel with = the prime science observations. The observations were executed as = scheduled, and no anomalies were noted. 1.2 Completed Four Sets of WF/PC-2 6188 (WF/PC-2 Cycle 5 Calibration: = Dark) The WF/PC-2 was used to make internal dark calibration images. The = observations were completed as scheduled, and no problems were reported. 1.3 Completed WF/PC-2 5972 (Determination of the Extragalactic Distance = Scale) The WF/PC-2 was used to make observations of the galaxy NGC-2090 = looking for standard candle Cepheid variable stars. The purpose of this = observations is to help estimate the Hubble constant with an accuracy of = 10%. The observations were executed as scheduled, and no anomalies were = noted. 1.4 Completed WF/PC-2 6250 (WF/PC-2 Cycle 5 Calibration: Internal = Monitor) The WF/PC-2 was used to make a series of internal calibration = observations. Internal flats, biases, and K-Spot calibration images were = taken. These observations are designed to monitor the health of the = cameras. The observations were executed as scheduled, and no anomalies = were noted. 1.5 Completed WF/PC-2 5989 (Imaging of CL1358+62 and Its Environs: the = Morphology-Density Relation)) The WF/PC-2 was used to make imaging observations of galactic = cluster CL1358+62 (aka, Gal-Clus-1) in order to study the morphology of = the galaxy as a function of its position in the cluster. The observations = were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 There was a loss of lock at 045/003938Z, after a successful = acquisition at 045/003352Z. The second acquisition attempt at 045/004320Z = was successful, but two of the WF-PC-2 observations for proposal 5657 were = lost when the take data flag was not up. These events are documented in = HSTAR 5502. Scheduled Reacquisitions : 4 Successful : 4 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Operations Notes: HST operations transferred to the SAMS backup facility for the = duration of the pre-simulation set-up and for the continuation and = completion of GSFC Simulation #1. No problems were encountered. ESTR #2 remains powered off pending conclusion of the high motor = current anomaly investigation. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH